For a special order the abc company manufactured 400 shirts sweatshirts were priced at $35 each and are $25 each The company received $11500 for the shirts How many of each type of shirt did the abc ompany manufacture for this order?
Found 2 solutions by Paul, sfischer18:
Answer by Paul(988)

Let $35 be x and let $25 be y
x+y=400
y=400-x

35x+25y=11500
35x+25(400-x)=11500
35x+10000-25x=11500
10x=1500
x=150
y=400-(150)
y=250

Hence, for $35 ---> 150 shirts and for $25 ---->250 shirts

Answer by sfischer18(2)

let x = $35
let y = $25
x+y=400
y=400-x
35x+25y=11500
35x+25(400-x)=11500
35x+10000-25x=11500
10x=1500
x=150
y=400-(150)
y=250
(150)= 150 sweatshirts
(250)= 250 tshirts